OpenStudy (anonymous):

Sixteen more than three-fifths of a number is two less five-ninths of the number. What is the number

OpenStudy (anonymous):

Let a be the number. From the statement of the problem: \[(3/5)* a+16=(5/9)* a-2\] \[a=-405\] \[(-405)*3/5+16=-227\] \[5/9*(-405)-2=-227 \]